A+B=140
B+C=160
C+D=130
D+A=110

Answer by tommyt3rd(5050)

Writing in matrix form we get the augmented matrix:

After row operations the system can be transformed to:

using t we can write:


A=110-t
B=t+30
C=130-t
D=t